# Aeos: Transport of metals from minihalos following Population III stellar feedback

### v1

@@ 4.1.2 Transport of Metals From the Halo
Figure 7 shows the evolution of the metal loss fraction $\ell$ with $M_{\rm dm}$. All four halos show the same characteristic shape in the evolution of $\ell$, albeit with variations, such as the metal injection event in Halo 0 at a mass of $\sim 3\times 10^{6}\,\rm M_{\odot}$, that likely depend on the time spacing of the chosen simulation outputs. Using more tightly spaced simulation outputs would likely reveal more of these variations. Figure 7 shows that following the first SNe, halos lose nearly all of the metals produced by stars within their virial radius. All four halos remain in this state of metal loss until they attain a mass of $\sim 10^{7}\,\rm M_{\odot}$. These halos then enter a period of rapid accretion whereby nearly all the metal mass produced returns to the halo (although Halo 2 does not explicitly show this, the last time step in our simulation shows a slight downturn that suggests it may be entering this stage). It is only at these high halo masses that halos no longer experience significant loss of the metals (or gas, see Figure 6) because of the injection of energy from new SNe.
We use Figure 8 to demonstrate that the reaccretion of metals is in fact the dominant contributor to the turnover in the metal loss fraction evolution. Figure 8 shows the evolution of the metal mass profiles for Halos 0 and 2, chosen for the differences in their evolution. If there were no evolution interior to the virial radius, and metal mass only increased via the smooth accretion of metals from the IGM, we would expect a smooth increase in metal mass from the outside-in over time. However, a number of distinct features are apparent. In both halos, a sudden increase in metal mass near the center of the halo is indicative of a SN injecting metals into the ISM. Subsequently, the metal mass profile relaxes, with metals in the center moving towards and beyond the outskirts of the halo. However, as the halo continues to evolve, the metal mass begins to build up again, from the outside-in, indicating accretion of metals. Thus, the reaccretion of metals is a significant contributor to the total amount of metals in each halo. While Halo 2 shows fairly simple evolution via smooth accretion following the first SNe, Halo 0 has a more complex evolution. Near the final redshift (yellow lines), we see that SNe once again evacuate metals from the center of the halo. However, in this case, the metals are not removed beyond the virial radius, an indicator that the potential well has grown deep enough for the halo to retain its metals.
@@
Notably, our results are similar to those of Kitayama & Yoshida (2005), who demonstrated that in low-mass halos ($M_{\rm dm}<10^{6}\,\rm M_{\odot}$), SN explosions, even with low input energy, lead to the complete evacuation of a halo’s gas in the presence of radiative feedback that lowered gas densities. Meanwhile, they found that in higher mass halos ($M_{\rm dm}\sim 10^{7}\,\rm M_{\odot}$), gas densities remain high and confine the expansion of the shock. Muratov et al. (2013) conducted an analysis similar to our study, and found that 80% of metals were retained within the virial radius for halos with $M_{\rm dm}>3\times 10^{6}\,\rm M_{\odot}$ at the first pair-instability SN event. We note that their analysis is not exactly comparable as the Muratov et al. (2013) simulations only resolve star formation in halos with $M_{\rm dm}>10^{6}\,\rm M_{\odot}$, whereas Aeos is able to resolve star formation in minihalos with $M_{\rm dm}>10^{5}\,\rm M_{\odot}$, implying that nearly all halos in Aeos that form stars have their first SNe at a mass of $M_{\rm dm}<3\times 10^{6}\,\rm M_{\odot}$ (see Figure 4).
@@ 5.1 Nucleosynthetic Contributions to Individual Elements
Given the yield table for Aeos and the set of stellar remnants that have produced Pop III or Pop II CCSNe, or have gone through an AGB phase, we can sum up the total injected metal mass as well as the injected metal mass of each individual element that each nucleosynthetic channel has contributed to their host halo. Figure 11 shows the breakdown of metal mass (total and individual elements) contributed by each channel to the four most massive halos in Aeos. The top row shows the cumulative star formation history broken down by Pop III and Pop II stars for each halo, enabling a comparison of the total stellar mass to the contribution in each channel. The remaining rows demonstrate that for each halo with a Pop II CCSNe contribution, the nucleosynthetic yields from Pop II quickly dominate the Pop III yields once Pop II CCSNe begin. However, with the exception of the neutron-capture elements (Sr and Ba), Pop II CCSNe dominate the nucleosynthetic production by only one or two orders of magnitude, despite the nearly three orders of magnitude dominance in mass for Pop II stars, due to the mass contribution from low mass stars. Figure 11 also highlights the wide diversity in star formation and enrichment histories seen across even just the four most massive halos, ranging from little to no impact from AGB stars, or even no component from Pop II stars at all.
We also find that AGB stars do not contribute much to early metal production due to the longer lifetimes of low mass stars. Even so, as the primary channel for s-process element production, their contribution to Sr and Ba is significant compared to other lighter elements. Though AGB stars contribute as much to Sr and Ba as Pop II CCSNe in our yield table (see Appendix of B24), we expect that at later times in the simulation, the AGB contribution to heavy elements will exceed the Pop II CCSN component as the result of a changed initial mass distribution of the available stars.
@@ 5.2 Loss Fraction of Individual Elements
Examining the representative elements shown in Figure 12, they share the same distinct features as the overall $\ell$ for all metals. This is true for all the elements that we follow, including those not shown. Specifically, we find that $\sim$100of individual metals are lost beyond the virial radius of the halo until the halo reaches a mass of $\sim 10^{7}\,\rm M_{\odot}$, and that this does not vary between elements from different nucleosynthetic families, including s-process elements produced by AGB stars. This is unsurprising given Figure 11, which shows that although AGB contributions to Sr and Ba are significant in comparison to their contributions to other elements, the contribution from CCSNe still dominates the production of Sr and Ba at these times. Thus, it is expected that we should see the same behavior for all elements at early times.
@@ 6 Conclusions
    Only $\lesssim 8$% of metal mass resides in halos until $z\approx 15.3$, at which point the most massive halos begin to retain metals. Those halos retain $\sim$92of all metals contained within halos (31–35% of total metals) at the end of our simulation at $z=14.3$.
@@
    Pop III star formation begins in halos with masses ranging from $10^{5}$ to $2\times 10^{6}\,\rm M_{\odot}$. At early times, these values are only a factor of 2–3 above the critical mass $M_{\rm crit}$ but within the 0.2 dex scatter derived by Kulkarni et al. (2021) for streaming velocity $v_{\rm bc}=0$ with a UV background strength of $J_{\rm LW}\sim 0-1\,J_{21}$.
@@
    Halos lose all gas and metals to beyond $r_{\rm{vir}}$ after SN enrichment events occur until they reach a dark matter mass of about $3\times 10^{6}$ and $1\times 10^{7}\,\rm M_{\odot}$ for gas and metals, respectively. The separate characteristic masses for gas and metals suggests that metals in hot shocked ejecta do not efficiently mix with the cooler local gas into which they are injected.
@@
Through its star-by-star treatment of star formation, Aeos is poised to study the chemical abundance patterns that result from the mixing of mass-dependent SN yields in both the ISM and IGM. The imprints of these patterns on the first enriched Pop II stars are observable, and the implications of these signatures for the conditions of the early Universe will fall to star-by-star cosmological simulations to interpret. Going forward, Aeos will be able to contribute to answering questions such as: what is the metal floor for enriched star formation (B24), on what scale do SN yields mix in the early Universe, how many SN abundance patterns are represented in the abundances of the first enriched stars, and how do enriched stellar abundances deviate from the local gas abundances and the predictions of homogeneous mixing models.
@@ Acknowledgments
J.M. acknowledges support from the NSF Graduate Research Fellowship Program through grant DGE-2036197. K.B. is supported by an NSF Astronomy and Astrophysics Postdoctoral Fellowship under award AST-2303858. G.L.B. acknowledges support from the NSF (AST-2108470 and AST-2307419, ACCESS), a NASA TCAN award, and the Simons Foundation through the Learning the Universe Collaboration. M.-M.M.L. and E.P.A. acknowledge partial support from NSF grant AST-2307950. A.P.J. acknowledges support by NSF grants AST-2206264 and AST-2307599. J.H.W. acknowledges support by NSF grant AST-2108020 and NASA grant 80NSSC21K1053. E.P.A. also acknowledges support from NASA ATP grant 80NSSC24K0935. A.F. acknowledges support from NSF grant AST-2307436.

### v2

@@ 4.1.2 Transport of Metals from the Halo
Figure 7 shows the evolution of the metal loss fraction $\ell$ with $M_{\rm dm}$. All four halos show the same characteristic shape in the evolution of $\ell$, albeit with variations, such as the metal injection event in Halo 0 at a mass of $\sim 3\times 10^{6}\,\rm M_{\odot}$, that likely depend on the time spacing of the chosen simulation outputs. Using more tightly spaced simulation outputs would likely reveal more of these variations. Figure 7 shows that, following the first SNe, halos lose nearly all of the metals produced by stars within their virial radius. All four halos remain in this state of metal loss until they attain a mass of $\sim 10^{7}\,\rm M_{\odot}$. These halos then enter a period of rapid accretion whereby nearly all the metal mass produced returns to the halo (although Halo 2 does not explicitly show this, the last time step in our simulation shows a slight downturn that suggests it may be entering this stage). It is only at these high halo masses that halos no longer experience significant loss of the metals (or gas, see Figure 6) because of the injection of energy from new SNe.
We use Figure 8 to demonstrate that the reaccretion of metals is in fact the dominant contributor to the turnover in the metal loss fraction evolution. Figure 8 shows the evolution of the metal mass profiles for Halos 0 and 2, chosen for the differences in their evolution. If there were no evolution interior to the virial radius, and metal mass only increased via the smooth accretion of metals from the IGM, we would expect a smooth increase in metal mass from the outside in over time. However, a number of distinct features are apparent. In both halos, a sudden increase in metal mass near the center of the halo is indicative of an SN injecting metals into the ISM. Subsequently, the metal mass profile relaxes, with metals in the center moving toward and beyond the outskirts of the halo. However, as the halo continues to evolve, the metal mass begins to build up again, from the outside in, indicating accretion of metals. Thus, the reaccretion of metals is a significant contributor to the total amount of metals in each halo. While Halo 2 shows fairly simple evolution via smooth accretion following the first SNe, Halo 0 has a more complex evolution. Near the final redshift (yellow lines), we see that SNe once again evacuate metals from the center of the halo. However, in this case, the metals are not removed beyond the virial radius, an indicator that the potential well has grown deep enough for the halo to retain its metals.
@@
Notably, our results are similar to those of Kitayama & Yoshida (2005), who demonstrated that, in low-mass halos ($M_{\rm dm}<10^{6}\,\rm M_{\odot}$), SN explosions, even with low input energy, lead to the complete evacuation of a halo’s gas in the presence of radiative feedback that lowered gas densities. Meanwhile, they found that, in higher-mass halos ($M_{\rm dm}\sim 10^{7}\,\rm M_{\odot}$), gas densities remain high and confine the expansion of the shock. Hicks et al. (2021) also found that halos $M_{\rm dm}<10^{6}\,\rm M_{\odot}$ do not retain their metals, and in fact, most of the metals eventually incorporated in star formation originated beyond the virial radius of the halo. Muratov et al. (2013) conducted an analysis similar to our study, and found that 80% of metals were retained within the virial radius for halos with $M_{\rm dm}>3\times 10^{6}\,\rm M_{\odot}$ at the first pair-instability SN event. We note that their analysis is not exactly comparable as the Muratov et al. (2013) simulations only resolve star formation in halos with $M_{\rm dm}>10^{6}\,\rm M_{\odot}$, whereas Aeos is able to resolve star formation in minihalos with $M_{\rm dm}>10^{5}\,\rm M_{\odot}$, implying that nearly all halos in Aeos that form stars have their first SNe at a mass of $M_{\rm dm}<3\times 10^{6}\,\rm M_{\odot}$ (see Figure 4).
@@ 5.1 Nucleosynthetic Contributions to Individual Elements
Given the yield table for Aeos and the set of stellar remnants that have produced Pop III or Pop II CCSNe, or have gone through an AGB phase, we can sum up the total injected metal mass as well as the injected metal mass of each individual element that each nucleosynthetic channel has contributed to their host halo. Figure 11 shows the breakdown of metal mass (total and individual elements) contributed by each channel to the four most massive halos in Aeos. The top row shows the cumulative star formation history broken down by Pop III and Pop II stars for each halo, enabling a comparison of the total stellar mass to the contribution in each channel. The remaining rows demonstrate that, for each halo with a Pop II CCSNe contribution, the nucleosynthetic yields from Pop II quickly dominate the Pop III yields once Pop II CCSNe begin. However, with the exception of the neutron-capture elements (Sr and Ba), Pop II CCSNe dominate the nucleosynthetic production by only 1 or 2 orders of magnitude, despite the nearly 3 orders of magnitude dominance in mass for Pop II stars, due to the mass contribution from low-mass stars. Figure 11 also highlights the wide diversity in star formation and enrichment histories seen across even just the four most massive halos, ranging from little to no impact from AGB stars, or even no component from Pop II stars at all.
We also find that AGB stars do not contribute much to early metal production due to the longer lifetimes of low-mass stars. Even so, as the primary channel for s-process element production, their contribution to Sr and Ba is significant compared to other lighter elements. Although AGB stars contribute as much to Sr and Ba as Pop II CCSNe in our yield table (see Appendix of B24), we expect that, at later times in the simulation, the AGB contribution to heavy elements will exceed the Pop II CCSN component as the result of a changed initial mass distribution of the available stars.
@@ 5.2 Loss Fraction of Individual Elements
Examining the representative elements shown in Figure 12, they share the same distinct features as the overall $\ell$ for all metals. This is true for all the elements that we follow, including those not shown. Specifically, we find that $\sim$100% of individual metals are lost beyond the virial radius of the halo until the halo reaches a mass of $\sim 10^{7}\,\rm M_{\odot}$, and that this does not vary between elements from different nucleosynthetic families, including s-process elements produced by AGB stars. This is unsurprising given Figure 11, which shows that, although AGB contributions to Sr and Ba are significant in comparison to their contributions to other elements, the contribution from CCSNe still dominates the production of Sr and Ba at these times. Thus, it is expected that we should see the same behavior for all elements at early times.
@@ 6.1 Summary
    Only $\lesssim 8$% of metal mass resides in halos until $z\approx 15.3$, at which point the most massive halos begin to retain metals. Those halos retain $\sim$92% of all metals contained within halos (31%–35% of total metals) at the end of our simulation at $z=14.3$.
@@
    Pop III star formation begins in halos with masses ranging from $10^{5}$ to $2\times 10^{6}\,\rm M_{\odot}$. At early times, these values are only a factor of 2–3 above the critical mass $M_{\rm crit}$ but within the 0.2 dex scatter derived by Kulkarni et al. (2021) for the streaming velocity $v_{\rm bc}=0$ with a UV background strength of $J_{\rm LW}\sim 0-1\,J_{21}$.
@@
    Halos lose all gas and metals to beyond $r_{\rm{vir}}$ after SN enrichment events, regardless of the number of SNe that occur, until they reach a dark matter mass of about $3\times 10^{6}$ and $1\times 10^{7}\,\rm M_{\odot}$ for gas and metals, respectively. The separate characteristic masses for gas and metals suggest that metals in hot shocked ejecta do not efficiently mix with the cooler local gas into which they are injected.
@@
6.2 Robustness to Simulation Choices
As with all simulations, the Aeos simulations make several choices that potentially influence the robustness of our results. We detail the effect that some of these choices may have below.
  1. 1.
    Altering the IMF may affect quantities such as the relative contribution of Pop III CCSNe to the enrichment history of minihalos, the metal floor [X/H] following Pop III enrichment, and the number of star forming halos due to varying strength of local ionizing sources. However, due to the independence of the metal loss fraction from the number of SNe exploding in the virial radius of the halo, we do not expect the characteristic mass of metal retention to change so long as Pop III CCSNe are still prevalent among halos.
  2. 2.
    If pair-instability SNe were included in our simulations, we might expect that the characteristic masses for gas and metal retention would increase due to the typical injection energy of $10^{53}$ erg, 2 orders of magnitude stronger than a single CCSN. An instantaneous energy injection of this magnitude could serve to blow metals and gas out farther than the typical CCSN, thereby increasing the halo mass at which they are reaccreted (as seen by the analysis of the SNe closely clustered in time in Halo 2) or to completely unbind the metals and gas from the original halo. This could result in a greater number of externally enriched halos and a greater extent of enrichment in the IGM.
  3. 3.
[…]
6.3 Implications for Numerical and Semi-Analytic Models
@@
Through its star-by-star treatment of star formation, Aeos is poised to study the chemical abundance patterns that result from the mixing of mass-dependent SN yields in both the ISM and IGM. The imprints of these patterns on the first enriched Pop II stars are observable, and the implications of these signatures for the conditions of the early Universe will fall to star-by-star cosmological simulations to interpret. Going forward, Aeos will be able to contribute to answering questions such as what is the metal floor for enriched star formation (B24), on what scale do SN yields mix in the early Universe, how many SN abundance patterns are represented in the abundances of the first enriched stars, and how do enriched stellar abundances deviate from the local gas abundances and the predictions of homogeneous mixing models.
@@ Acknowledgments
The authors wish to thank the anonymous referee for a thorough review that has improved the clarity and completeness of the manuscript.
J.M. acknowledges support from the NSF Graduate Research Fellowship Program through grant DGE-2036197. K.B. is supported by an NSF Astronomy and Astrophysics Postdoctoral Fellowship under award AST-2303858. G.L.B. acknowledges support from the NSF (AST-2108470 and AST-2307419, ACCESS), a NASA TCAN award, and the Simons Foundation through the Learning the Universe Collaboration. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). M.-M.M.L. and E.P.A. acknowledge partial support from NSF grant AST-2307950. A.P.J. acknowledges support by NSF grants AST-2206264 and AST-2307599. J.H.W. acknowledges support by NSF grant AST-2108020 and NASA grant 80NSSC21K1053. E.P.A. also acknowledges support from NASA ATP grant 80NSSC24K0935. A.F. acknowledges support from NSF grant AST-2307436.